OpenStudy (anonymous):

find the derivative and all non zero x-coordinates at which the function gx=x^(a)*(e^(Bx)) has horizontal tagent lines using using product and chain rules and logarithmic differentiation

OpenStudy (anonymous):

You have to use both methods mentioned above and the function looks like this: \[f(x)=x ^{\alpha}e ^{\beta x}\]

OpenStudy (anonymous):

\[g(x)=x^ae^{bx}\] \[g'(x)=ax^{a-1}e^{bx}+x^ae^{bx}\times b\] by the chain rule

OpenStudy (anonymous):

or if you prefer \[\large g(x)=\alpha x^{\alpha -1}e^{\beta x}+\beta x^{\alpha}e^{\beta x}\]

OpenStudy (anonymous):

factor out the common factor of \(e^{\beta x}\) set the rest equal to zero, as \(e^{\beta x}\) is never zero

OpenStudy (anonymous):

Okay, let me give that a shot!

OpenStudy (anonymous):

Um, question? My answer key says the derivative is \[x ^{\alpha -1}e ^{\beta x} (\alpha+\beta x)\] any idea how they got there?

OpenStudy (anonymous):

Hello? Um, help, please?